Juan mixes 25 mL of a 15% acid solution with 60 mL of a 25% acid solution and 48 mL of a 50% acid solution. About what percent acid is the final solution?
If you could please help me figure this out with using only one variable (perferrably x) that would be great...thanks!!!

Answer by mananth(16946)

Juan mixes 25 mL of a 15% acid solution with 60 mL of a 25% acid solution and 48 mL of a 50% acid solution. About what percent acid is the final solution?
25ml 15%
60 ml 25%
48 ml 50%
Let percent in thefinal solution be x
and quantity will be 133 ml
25*15+60*25+48*50=133*x
375+1500+2400=133x
4275 =133x
x= 424/133
x= 32.14%